# CRVQ: Channel-relaxed Vector Quantization for Extreme Compression of LLMs

### v1

@@ 4.2 Channel Reordering
Although the importance of all channels is computed, we are not yet able to handle the more critical channels independently by VQ framework. This is because the distribution of important channels is highly sparse, and their neighboring channels are often not critical. In other words, critical channels are not inherently contiguous. To break this obstacle, we reorder all channels based on their importance. As shown in the left of Figure 3, the scattered distribution of important channels in the weight matrix is reorganized such that all critical channels become adjacent.
4.3 Multi-codebook Fitting
@@
We observe that channel importance varies sharply, with a very small proportion of channels being highly critical. Based on this observation, we conceive that applying higher-precision quantization to a few critical channels could lead to substantial performance improvements at minimal additional cost. In VQ, there are two primary paths to enhance quantization precision, using more codebooks or employing codebooks with larger capacity. However, increasing codebook capacity also increases the encoding bit-width $e$, which in turn inflates the code for non-critical vectors. Therefore, we opt to use additional codebooks instead of expanding the capacity of existing ones. In CRVQ, we designate the codebooks applied to all vector groups as basic codebooks and those specifically used for critical vector groups as extended codebooks. As shown in the middle of Figure 3, codebook 1 is the basic codebook, while codebooks 2 to 4 are extended codebooks. We first perform quantization on all groups using the basic codebook. Next, we supplement the quantization of critical vector groups by fitting them with the extended codebooks using additive VQ. Additive quantization iteratively fits the residual from the previous quantization step, allowing critical vector groups to benefit from multiple rounds of refinement. This multi-codebook fitting enables critical vector groups to work with higher precision, thereby playing a more important role in maintaining overall performance. During inference, the weights are restored to their original order through a reverse reordering process.
The additive quantization employed in our method is based on the AQLM algorithm, which utilizes multi-round fine-tuning and beam search to enhance the precision of additive quantization [14]. In our CRVQ, we quantize the model layer by layer and end-to-end fine-tuning is also introduced. A key distinction from AQLM, which applies multiple codebooks to all vector groups, lies in our approach: we identify and isolate a small subset of critical channels and selectively apply multi-codebook only to these channels. Please refer to Algorithm 1 for details.
@@
8:         $\mathbf{W}_{\text{encoded}},\mathcal{B}_{\text{base}}\leftarrow\texttt{VectorQuantize}(\mathbf{W}_{\text{sorted}},\mathcal{C}_{\text{base}})$triangleright $\mathcal{B}$ is the encoding of $\mathbf{W}$ on codebook $\mathcal{C}$
@@
12:             $\mathbf{E}^{t}_{\text{encoded}},\mathcal{B}_{\text{ext}}^{t}\leftarrow\texttt{VectorQuantize}(\mathbf{E}^{t},\mathcal{C}_{\text{ext}}^{t})$
13:             $\mathbf{W}_{\text{encoded}}^{\lambda}\leftarrow\texttt{CombineWeights}(\mathbf{W}_{\text{encoded}}^{\lambda},\mathbf{E}^{t}_{\text{encoded}})$
@@
[…]
18:         end while
19:    end for
20:    $\texttt{FineTuneBlock}(B_{i})$
21:end for
22:return $\mathcal{M}_{\text{quant}}\leftarrow\texttt{E2E\_FineTune}(\mathcal{M})$
@@ 5.1 Settings
We perform experiments on OPT-125M/1.3B, LLaMA-7B/13B, and LLaMA2-7B/13B to evaluate our method. Since the baselines used for comparison also require calibration data, we follow the experimental setup of AQLM by randomly sampling fixed-length data from the Red_Pajama dataset to support algorithm execution. For LLaMA2 models, the data length is set to 4096, while for the other models, calibration is performed with a text length of 2048.
Baselines
We compare CRVQ with several representative extreme PTQ methods, including PB-LLM [40], BiLLM [21], and AQLM [14]. PB-LLM can be integrated with other classical PTQ algorithms. The reported results for PB-LLM use GPTQ as its backbone, with 10% of critical weights quantized to 8 bits to ensure a high compression ratio, achieving an average bit-width below 2 bits. BiLLM employs weight partition and residual-based multi-step quantization to improve accuracy. Following the original paper, we use a block size of 128 and 128 calibration samples. AQLM is a vector quantization method that combines layer-level, block-level, and end-to-end fine-tuning. To achieve 1-bit quantization, we configure the codebook count $m=1$, vector dimension $d=8$, and codebook bit-width $e=8$. We adopt its optimal setup by using 2048 calibration samples. For our proposed CRVQ, the ratio of critical channels is set to $\lambda=2\%$, with one basic codebook and three extended codebooks. Similar to AQLM, CRVQ employs parameters of $d=8$, $e=8$, and 2048 calibration samples. All experiments are conducted on 1$\times$×NVIDIA A100-80GB. Further details on the baselines are provided in Section A.1.
Evaluation Metrics
To evaluate the performance of various baselines, we calculate perplexity using datasets like WikiText2 [27] and C4 [28]. The evaluation data is randomly sampled and its length matches that of the corresponding calibration data. A lower perplexity score reflects better preservation of the original output distribution. Moreover, we report the zero-shot accuracies on commonsense reasoning tasks such as Winogrande [29], HellaSwag [41], PIQA [4], BoolQ [7], and ARC-e/ARC-c [8]. These tasks measure the extent to which the quantized model retains the original model capability on downstream applications. For these evaluations, we leverage the open-source toolkit “LM-Evaluation-Harness” [^1] https://github.com/EleutherAI/lm-evaluation-harness to conduct zero-shot task assessments.
Method Wbits Perplexity($\downarrow$) Zero-shot Accuracy($\uparrow$)
@@
OPT-125M
PB-LLM 1.73 4235.91 3031.30 49.04 24.97 49.82 37.80 25.02 20.16 34.47
BiLLM 1.11 3818.77 2502.60 50.67 25.74 50.92 37.83 26.47 24.74 36.06
AQLM 1.08 1054.03 828.22 50.67 26.96 51.74 37.83 27.06 23.81 36.34
CRVQ 1.27 641.21 533.07 49.57 27.99 52.07 37.83 28.37 22.78 36.44
OPT-1.3B
PB-LLM 1.73 912.41 834.57 50.27 27.14 52.10 38.10 29.46 22.51 36.59
BiLLM 1.11 1099.54 707.30 49.96 27.32 53.54 37.80 29.29 20.73 36.44
AQLM 1.02 809.28 674.63 50.28 27.14 50.38 37.83 27.61 24.15 36.23
CRVQ 1.10 303.24 266.84 50.28 28.03 52.29 45.54 30.30 22.61 38.18
LLaMA-7B
PB-LLM 1.71 231.79 272.66 50.38 28.89 53.26 44.55 31.82 23.17 38.67
@@
5.2 Main Results
Table 1 lists the results of strong baselines as well as CRVQ on different models. In various model sizes, CRVQ clearly outperforms the baselines. First, although PB-LLM has the highest average bit-width compared to other baselines, it shows no clear advantage in perplexity or zero-shot tasks. This may be attributed to the suboptimal bit utilization in earlier methods. In contrast, BiLLM, AQLM, and CRVQ show comparable performance. Moreover, we observe that although extreme PTQ incurs certain performance loss compared to QAT [37, 23], its benefits grow rapidly with increasing model size. For example, while quantized models such as those based on smaller OPT, exhibit noticeable performance degradation, the algorithms perform increasingly better on larger LLaMA models, reaching performance levels similar to QAT [37] or even unquantized models [33].
For perplexity, our method achieves the best results across all models and all competing methods, with AQLM being the second-best. For instance, our method achieves a perplexity of 9.81 on the Wiki2 dataset with LLaMA-13B, while AQLM achieves 15.25. This demonstrates the remarkable capability of VQ in reducing weight bit-width. More importantly, the negligible additional bit-width cost of CRVQ relative to AQLM highlights the significance and effectiveness of applying multi-codebook fitting to critical channels.
For zero-shot accuracy, our method achieves optimal results in most cases. Notably, CRVQ demonstrates the most noticeable advantages within the LLaMA series, apart from a few tasks on LLaMA-13B, such as Winogrande and ARC-e. For most LLaMA models, our approach provides performance improvements of 7% to 13% compared to the second-best baseline.
6 Discussion
@@ 6.1 Ratio of Important Channels
The core idea of CRVQ is to leverage additional codebooks for a few critical channels, achieving maximum quantization benefits with minimal overhead. A natural question arises: how many channels are sufficient for CRVQ to perform effectively? We conduct experiment under codebook count $m=2$ $(8+8\times 1)$, varying the channel ratio $\lambda$ from 5e-3 to 0.1. As shown in Figure 4(a), when $\lambda<0.02$, the perplexity of the quantized model decreases rapidly, leading a reduction of 6.7 compared to the model without critical channels. Increasing $\lambda$ to 0.1 further improves performance, even though the average bit-width remains as low as 1.1 bit. These results validate the significant impact of critical channels. Considering the performance-to-cost ratio at $\lambda=0.02$, this setting maybe well-suited for deployment on extremely low-resource devices.
@@ 6.2 Number of Extended Codebooks
In CRVQ, we apply multiple extended codebooks to a few critical channels. There are also two important questions: how many extended codebooks are necessary for the method to work effectively, and is more always better? We investigate the impact of the extended codebooks on performance under two importance ratios, $\lambda=0.02$ and $\lambda=0.1$. As shown clearly in Figure 4(b) and 4(c), using three extended codebooks provides the most notably performance improvement compared to using none in both settings. However, when the extended codebooks count exceeds three, the performance gains become marginal. Thus, adding more codebooks does not necessarily yield better results. Based on our findings, we recommend using three extended codebooks for optimal performance.
6.3 Different Reorder Strategy
We conduct experiments and analyses on different importance metrics to understand the varying contributions of different channels in VQ. Based on the analysis in Section 4.1, we evaluate three channel reordering strategies: random reordering (Random), reordering based on weight magnitude (W-only), and reordering using weight metric combined with Hessian proxy (W-A). The results presented in Table 2 indicate that the W-A consistently outperforms the alternatives. We also observed that the effectiveness of Random and W-only strategies varies with $\lambda$. Specifically, W-only performs better at extremely low $\lambda$. It is likely because the Random strategy has low hit rates for critical channels under such conditions. Overall, our findings highlight the necessity of explicitly catching critical channels in VQ to achieve superior performance.
Table 2: Performance on different reorderding strategy. We use $8+8\times 3$ setting.
Reorder Strategy Perplexity($\downarrow$) Zero-shot Accuracy($\uparrow$)
@@ 6.3 Different Reorder Strategy
6.4 Combination with Stronger VQ Settings
@@
Our proposed CRVQ is not limited to settings with a single base codebook and codebook bit-width of $e=8$. On the contrary, it can be seamlessly integrated with any vanilla VQ method, applying multi-codebook fitting specially to critical channels. First, as the bit-width of the base codebook increases, its fitting capability improves, but CRVQ still benefits from applying additional codebooks to a few critical channels. Figure 5 presents the performance of CRVQ when applied to two larger codebooks (e = 10 and e = 12). It can be observed that CRVQ consistently shows effectiveness as the extended codebooks increase. Moreover, our method also works well with involving more base codebooks. For instance, under the setting of $m=2,d=16,e=8$, the perplexity of AQLM is 17.11, whereas CRVQ achieves 14.21 with three extended codebooks. This demonstrates the generality and flexibility of our approach across diverse settings.

### v2

@@ 4.2 Channel Reordering
Although the importance of all channels is computed, we are not yet able to handle the more critical channels independently by VQ framework. This is because the distribution of important channels is highly sparse, and their neighboring channels are often not critical. In other words, critical channels are not inherently contiguous. To break this obstacle, we reorder all channels based on their importance. As shown in the left of Figure 3, the scattered distribution of important channels (index 1,3 and 5) in the weight matrix is reorganized such that all important channels become adjacent (in the left vector group).
4.3 Extended Codebook Fitting
@@
There are two primary paths to enhance quantization precision, using more codebooks or employing codebooks with larger capacity. However, increasing codebook capacity also increases the encoding bit-width $e$, which in turn inflates the code for non-critical vectors. Therefore, we opt to use additional codebooks instead of expanding the capacity of existing ones. In CRVQ, we designate the codebooks applied to all vector groups as basic codebooks and those specifically used for critical vector groups as extended codebooks. As shown in the middle of Figure 3, codebook 1 is the basic codebook, while codebooks 2 to 4 are extended codebooks. We first perform quantization on all groups using the basic codebook. Next, we supplement the quantization of critical vector groups by fitting them with the extended codebooks using additive VQ. During inference, the weights are restored to their original order through a reverse reordering process. In this paper, we use notation like “$8+8\times 3$” to denote the codebook configuration, where one basic codebook and three extended codebooks are used.
4.4 Overall Algorithm
Now we demonstrate our CRVQ with the help of Algorithm 1. Similar to most PTQ methods, the quantization process is performed layer-by-layer. For each weight matrix $\mathbf{W}$ in a specific layer, we first compute the column-wise channel importance and reorder these channels (Line $3\sim 5$). Then, we construct the basic codebook $\mathcal{C}_{\mathrm{base}}$ and search the binary encoding $\mathcal{B}_{\mathrm{base}}$ as well as the quantized weight $\mathbf{W}_{\mathrm{encoded}}$ (Line $6\sim 7$). Here we use only one basic codebook and please refer to Section 3.2 for more details. Next, we recur to $m-1$ extended codebooks $\{\mathcal{C}^{1}_{\text{ext}},\ldots,\mathcal{C}^{m-1}_{\text{ext}}\}$ to fitting the $\lambda$ important channels in additive manner (Line $8\sim 13$). Additive VQ iteratively fits the residual $\mathbf{E}^{t}$ from the previous quantization step, allowing important vector groups to benefit from multiple rounds of refinement. Finally, we follow previous works (Egiazarian et al., 2024; Tseng et al., 2024) by finetuning the codebook and beam-searching the optimal code (Line $14\sim 17$). Moreover, block and end-to-end finetuning is leveraged for enhancing the quantization performance (Line $19,21$), which is widely used in Egiazarian et al. (2024); Tseng et al. (2024).
@@ 4.4 Overall Algorithm
8:         $\mathbf{W}_{\text{encoded}},\mathcal{B}_{\text{base}}\leftarrow\texttt{VectorQuant}(\mathbf{W}_{\text{sorted}},\mathcal{C}_{\text{base}})$triangleright $\mathcal{B}$ is the encoding of $\mathbf{W}$ on codebook $\mathcal{C}$
@@
12:              $\mathbf{E}^{t}_{\text{encoded}},\mathcal{B}_{\text{ext}}^{t}\leftarrow\texttt{VectorQuant}(\mathbf{E}^{t},\mathcal{C}_{\text{ext}}^{t})$
13:              $\mathbf{W}_{\text{encoded}}^{\lambda}\leftarrow\texttt{Update}(\mathbf{W}_{\text{encoded}}^{\lambda},\mathbf{E}^{t}_{\text{encoded}})$
@@
15:         while $\texttt{QuantLoss}(\mathbf{W},\mathbf{W}_{\text{encoded}},\mathbf{X})\geq\epsilon$ do $\triangleright$ loss is $\|\mathbf{WX}-\mathbf{W}_{\text{encoded}}\mathbf{X}\|_{2}^{2}$
@@
18:     $\{\mathcal{C}^{1}_{\text{ext}},\ldots,\mathcal{C}^{m-1}_{\text{ext}}\},\mathcal{B}_{\text{base}},\{\mathcal{B}^{1}_{\text{ext}},\ldots,\mathcal{B}^{m-1}_{\text{ext}}\})$
19:         end while
20:     end for
21:     $\texttt{FineTuneBlock}(B_{i})$
22:end for
23:return $\mathcal{M}_{\text{quant}}\leftarrow\texttt{E2E\_FineTune}(\mathcal{M})$
@@ 5.1 Settings
We perform experiments on OPT-1.3B/2.7B, LLaMA-7B/13B and LLaMA2-7B/13B to evaluate our method. Results on LLaMA3 and Qwen2 are in Section A.3. Since the baselines used for comparison also require calibration data, we follow the experimental setup of AQLM by randomly sampling fixed-length data from the Red_Pajama dataset to support algorithm execution. For OPT models, the data length is set to 2048, while for the other models, calibration is performed with a text length of 4096.
Model Method Wbits Perplexity($\downarrow$) Zero-shot Accuracy($\uparrow$)
@@
OPT-1.3B FP16 16.0 14.67 14.76 59.67 53.69 72.25 56.85 51.43 29.86 53.96
[…]
BiLLM 1.11 59.24 67.71 52.33 32.87 58.00 52.17 35.52 23.04 42.32
AQLM 1.02 59.68 45.82 50.75 29.94 55.71 52.26 33.21 23.29 40.87
CRVQ 1.08 39.73 33.29 52.80 32.87 58.43 52.11 36.15 23.63 42.67
OPT-2.7B FP16 16.0 12.46 13.17 60.69 60.64 74.59 59.60 54.34 31.23 56.85
[…]
BiLLM 1.11 46.40 47.95 51.85 33.82 59.41 52.17 35.61 23.46 42.73
AQLM 1.01 38.73 32.51 50.43 31.61 57.56 51.77 35.61 23.81 41.80
CRVQ 1.08 28.91 26.11 52.88 36.09 60.94 51.99 39.35 25.09 44.40
LLaMA-7B FP16 16.0 5.68 7.08 69.93 76.16 79.16 74.98 72.90 44.62 69.62
[…]
@@
Baselines
We compare CRVQ with several representative extreme PTQ methods, including PB-LLM (Yuan et al., 2024), BiLLM (Huang et al., 2024), and AQLM (Egiazarian et al., 2024). PB-LLM is integrated with GPTQ as its backbone and 10% of critical weights quantized to 8 bits to ensure an average bit-width below 2 bits. In BiLLM, we use a block size of 128 and 128 calibration samples. In AQLM, we configure the codebook count $m=1$, vector dimension $d=8$, and codebook bit-width $e=8$. We adopt its optimal setup by using 2048 calibration samples. For our proposed CRVQ, the ratio of critical channels is set to $\lambda=2\%$, with one basic codebook and three extended codebooks. Similar to AQLM, CRVQ employs $d=8$, $e=8$, and 2048 calibration samples. All experiments are conducted on 1$\times$×NVIDIA A100-80GB. Further details on the baselines are provided in Section A.9.
Evaluation Metrics
To evaluate baseline performance, we calculate perplexity on datasets like WikiText2 (Merity et al., 2017) and C4 (Raffel et al., 2020), using randomly sampled evaluation data of the same length as the calibration data. Lower perplexity indicates better preservation of the original output distribution. Additionally, we report zero-shot accuracies on commonsense reasoning tasks, including Winogrande (Sakaguchi et al., 2021), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), BoolQ (Clark et al., 2019), and ARC-e/ARC-c (Clark et al., 2018). Evaluations are conducted using the open-source toolkit LM-Evaluation-Harness [^1] https://github.com/EleutherAI/lm-evaluation-harness.
@@
5.2 Main Results
As shown in Table 5.1, CRVQ consistently outperforms baselines across model sizes. While PB-LLM has the highest average bit-width, it shows no clear advantage in perplexity or zero-shot tasks, likely due to suboptimal bit utilization. In contrast, BiLLM, AQLM, and CRVQ perform comparably. Extreme PTQ incurs some performance loss compared to QAT (Xu et al., 2024; Jo et al., 2024), but its benefits scale with model size. Smaller models like OPT show performance degradation, whereas larger LLaMA models achieve results close to QAT (Xu et al., 2024) or even unquantized models (Touvron et al., 2023b).
For perplexity, CRVQ achieves the lowest across all models, with AQLM as the second-best. For example, on WikiText2 with LLaMA-13B, CRVQ achieves a perplexity of 9.81, compared to 15.25 for AQLM, demonstrating their effectiveness in reducing bit-width. More importantly, CRVQ achieves this with negligible additional bit-width cost.
For zero-shot accuracy, CRVQ achieves the best results in most cases, particularly within the LLaMA series, except for a few tasks on LLaMA-13B (e.g., Winogrande, ARC-e). On most LLaMA models, CRVQ outperforms the second-best baseline by 7% to 13%.
6 Discussion
@@ 6.1 Ratio of Important Channels
CRVQ leverages extra codebooks for a few critical channels to maximize quantization benefits with minimal overhead. A natural question arises: how many channels are sufficient for CRVQ to perform effectively? We test this under codebook count $m=2$, i.e. “$8+8\times 1$”, varying the channel ratio $\lambda$ from 5e-3 to 0.1. As shown in Figure 4(a), when $\lambda<0.02$, the perplexity of the quantized model drops rapidly by 6.7 compared to the model without critical channels. Increasing $\lambda$ to 0.1 further improves performance, even though the average bit-width remains as low as 1.1 bit. These results validate the significant impact of critical channels. Considering the performance-to-cost ratio at $\lambda=0.02$, this setting maybe well-suited for deployment on extremely low-resource devices.
@@ 6.2 Number of Extended Codebooks
In CRVQ, we apply multiple extended codebooks to a few critical channels. There are also two important questions: how many extended codebooks are necessary for the method to work effectively, and is more always better? We investigate the impact of the extended codebooks under two critical channel ratios, $\lambda=0.02$ and $\lambda=0.1$. As shown clearly in Figure 4(b) and 4(c), using three extended codebooks provides the most notably performance improvement compared to using none in both settings. However, when the extended codebooks count exceeds three, the performance gains become marginal. Thus, adding more codebooks does not necessarily yield better results and we recommend using three extended codebooks for optimal performance.
Ratio Reorder Strategy Perplexity($\downarrow$) Zero-shot Accuracy($\uparrow$)
@@
Table 2: Performance on different reorderding strategy. Here we use $8+8\times 3$ setting.
6.3 Different Reorder Strategy
We discuss on different importance metrics to understand the contributions of different channels in CRVQ. Based on the analysis in Section 4.1, we evaluate three channel reordering strategies: random reordering (Random), reordering based on weight magnitude (W-only), and reordering using weight combined with Hessian proxy (W-A). The results in Table 2 indicate that the W-A consistently outperforms the others. We also find that the effectiveness of Random and W-only strategies varies with $\lambda$. Specifically, W-only performs better than Random at extremely low $\lambda$. It is likely because the Random strategy has low hit rates for critical channels under such condition. These results emphasize the necessity of explicitly catching critical channels in VQ to achieve superior performance.
@@ 6.4 Flexible Codebook Settings
CRVQ is not limited to integrate with a single basic codebook and $e=8$. On the contrary, it can be seamlessly integrated with any vanilla VQ method, applying multi-codebook fitting specially to critical channels. First, as the basic codebook bit-width increases, its fitting capability improves, but CRVQ still works well. Figure 5 presents the performance of CRVQ when applied to two larger codebooks ($e$ = 10 and $e$ = 12). Obviously, CRVQ consistently shows effectiveness as the extended codebooks increase. Moreover, our method also works well with involving more base codebooks. For instance, under the setting of $m=2,d=16,e=8$, the perplexity of AQLM is 17.11, whereas CRVQ achieves 14.21 with three extended codebooks. This demonstrates the generality and flexibility of our approach across diverse settings.
6.5 Hardware-friendly Deployment